Mrs. Johnson invested $4000; part at 5%, the remainder at 3%. The total annual income from both investments was $152. Find the amount invested at each rate. Needs a legend?
Answer by lwsshak3(11628)

Mrs. Johnson invested $4000; part at 5%, the remainder at 3%. The total annual income from both investments was $152. Find the amount invested at each rate. Needs a legend?
let x= amt invested at 5%
4000-x= amt invested at 3%
5%x+3%(4000-x)=152
.05x+120-.03x=152
.02x=32
x=1600
4000-x=2400
amt invested at 5%=$1600
amt invested at 3%=$2400